OpenStudy (anonymous):

Given the function f(x)=x^3+x^2-2x+1, what is the resulting function when f(x) is shifted to the left one unit

ganeshie8 (ganeshie8):

simply replace each occurrence of `x` by `x+1`

terenzreignz (terenzreignz):

Still trying to find a way to explain as simply as possible WHY that action would do the desired shift.. :D oh well ,never mind

OpenStudy (anonymous):

So f(x+1)=(x+1)^3+(x+1)^2-2(x+1)+1?

ganeshie8 (ganeshie8):

bit tricky to explain - may be we need to go back in time and see how to fix the origin

terenzreignz (terenzreignz):

As I said.. never mind. ^^

ganeshie8 (ganeshie8):

Excellent ! @Majestic4321 thats right !!

OpenStudy (anonymous):

So then it would reduce to f(x-1)=x^3-2x^2-x+3?

OpenStudy (anonymous):

\[f(x)=(x+1)^{3}+ (x+1)^{2}-2(x+1)+1\]

OpenStudy (anonymous):

Right but that's not a choice so it would have to reduce right?

OpenStudy (anonymous):

Is this the choice

OpenStudy (anonymous):

No.name yes it is

OpenStudy (anonymous):

then that is indeed the choice ,

OpenStudy (anonymous):

Thank you:) I second guess myself a lot

OpenStudy (anonymous):

no problem

ganeshie8 (ganeshie8):

hey no, wait a sec

OpenStudy (anonymous):

What

ganeshie8 (ganeshie8):

` f(x-1)=x^3-2x^2-x+3` This is shift right transformation - Not the one the question was asking you to find

ganeshie8 (ganeshie8):

For left shift, ou need to find : f(x\(\color{Red}{+}\)1)

OpenStudy (anonymous):

Oh... SO than it would be f(x+1)=x^3+4x^2+3x+1?

Parth (parthkohli):

I think all of the choices are wrong... \(f(x + 1) = x^3 + 4x^2 + 3x + 1\) and \(f(x) = x^3 + x^2 - 2x + 1\) represent the same function. :P

ganeshie8 (ganeshie8):

yes they refer to the function f itself ! f(x+1)=x^3+4x^2+3x+1 is \(\large \color{red}{\checkmark}\) @Majestic4321

OpenStudy (anonymous):

My options with positive one are f(x)+1=x^3+x^2-2x+2 or f(x+1)=x^3+4x^2+3x+1

OpenStudy (anonymous):

Okay thanks haha

ganeshie8 (ganeshie8):

np, you seem new here... let me take a moment to welcome you :) \(\large \color{green}{\text{Welcome to Openstudy !}}\)

OpenStudy (anonymous):

Yes I am thank you:)